What is shape of phase space trajectory of classsical harmonic oscillator.
1
Expert's answer
2020-10-29T07:02:05-0400

The simplest Autonomous vibrating system is called "harmonic oscillator"; its dynamics is described by linear differential equation of the form:

"{\displaystyle {\ddot {x}}+\omega _{0}^{2}x}=0"

Such a system performs periodic sinusoidal (harmonic) movements; oscillatory motion does not occur only in the case "{\displaystyle x_{0}=0}" "x_{0}=0" and "{\displaystyle {\dot {x}}_{0}=0}" "{\displaystyle {\dot {x}}_{0}=0}" , that is, when the oscillator is initially in a state of equilibrium — in this case, it continues to remain in it. The coordinate equation of the phase trajectory of such a system defines integral curves in the form of a family of similar (with a constant ratio of axes) ellipses, with one and only one ellipse passing through each point of the phase path. This state of equilibrium is a special point of this system, namely, the center. Sourse:https://en.wikipedia.org/wiki/Phase_space